OpenStudy (anonymous):

Tristan says that solutions to radical equations are extraneous. Hadley says that solutions to radical equations are non-extraneous. Find an equation where Tristan is correct and second equation where Hadley is correct. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

not sure what the question is asking are you trying to find a radical equation with an extraneous solution?

OpenStudy (anonymous):

fiing to equations where there are both right

OpenStudy (anonymous):

here is a radical equation \[x=\sqrt{2x+3}\] if you square both sides you get \[x^2=2x+3\] or \[x^2-2x-3=0\] solutions are \(3\) and \(-1\) \(-1\) is extraneous, \(3\) is an actual solution

OpenStudy (anonymous):

are you going to call me an idiot again @me

OpenStudy (mertsj):

Tristan: \[\sqrt{x-2}=-3\] \[x-2=9\] \[x=11\] Check: \[\sqrt{11-2}=-3\] \[\sqrt{9}=-3\] \[3\neq-3, extraneous\]

OpenStudy (mertsj):

Hadley: \[\sqrt{x-2}=3\] \[x-2=9\] \[x=11\] Check: \[\sqrt{11-2}=3\] \[\sqrt{9}=3\] \[3=3, non-extraneous\]